OpenStudy (anonymous):

PLEASE HELP!!!! The probability of a student scoring 75% in class work is 0.64, and the probability of a student scoring 85% is 0.45. Event A: The student scores 75%. Event B: The student scores 85%. The probability of a student scoring 85% in class work, given that they have already scored 75% in class work, is 0.55. The probability of a student scoring 75% in class work, given that they have already scored 85% in class work, is 1. Which statement is true?

OpenStudy (anonymous):

Events A and B are independent because P(A|B) = P(A). Events A and B are independent because P(A|B) = P(B). Events A and B are independent because P(A|B) = P(A) + P(B). Events A and B are not independent because P(A| B) ≠ P(A). Events A and B are not independent because P(A|B) = P(A).

OpenStudy (amistre64):

how do we define P(A|B) ?

OpenStudy (anonymous):

I honestly do not know. I never learned this! I had a very horrible geometry teacher who caused me to fail due to lack of teaching! Im currently on plato making up the credit! this is my final exam!

OpenStudy (amistre64):

the best time to study for a test is before the test .... not during the test :/

OpenStudy (anonymous):

@amistre64 its probibality taken AND with probibality of B

OpenStudy (anonymous):

it never gave me anything to study, otherwise I would have!

OpenStudy (amistre64):

the test for independance is such that: